# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v4

@@ Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around key pillars: Explainability, ModelOps, Security, Privacy and their Lifecycle Governance, each contextualized to the challenges of AMAS. A risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows. We further discuss strategies for improving explainability in Agentic AI, as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, highlighting key directions to align emerging systems with TRiSM principles-ensuring safety, transparency, and accountability in their operation.
@@ 1 Introduction
AI governance has moved from aspiration to obligation. The EU Artificial Intelligence Act [1] entered into force on 1 August 2024 with a staged application (e.g., prohibitions in early 2025 and obligations for general-purpose AI from 2 August 2025  [^3] EU AI Act); meanwhile, organizations are adopting management frameworks such as ISO/IEC 42001:2023 [2] and the NIST AI Risk Management Framework (AI RMF 1.0) [3] to operationalize risk controls. In parallel, enterprise use of AI continues to expand: recent global surveys [4, 5, 6] report that a substantial majority of organizations use AI in at least one business function. These trends sharpen demand for system-level trust, risk, and security management tailored to agentic systems.
Figure 1: TRiSM taxonomy for Agentic AI (AMAS) presented in this review. We cover: (i) Fundamentals of Agentic AI (Agentic vs. traditional agents; LLM-based architectures), (ii) AI Trust, Risk, and Security Management (TRiSM) pillars: Explainability, ModelOps, Security, Privacy, and Lifecycle Governance, (iii) Trust management via explainability, (iv) Threats and risks, and (v) Evaluation metrics, including human-centered and composite measures.
@@
An AI agent is a computational entity that perceives its environment and acts to achieve goals [7]. What were once task-specific and largely deterministic programs have rapidly evolved into LLM-powered agentic systems with planning, tool use, and persistent memory [8]. In this paper, we refer to such LLM-based, coordinating agent systems as Agentic multi-agent systems (AMAS). These systems orchestrate multiple specialized agents to address long-horizon tasks, coordinate functions and roles, and adapt workflows through interactions with tools, users, and other agents. The shift from single-agent or rule-based pipelines to AMAS unlocks new capabilities but complicates coordination, control, transparency, and risk management across heterogeneous components and contexts.
Why TRiSM for Agentic AI?
@@
LLM-based AMAS exhibit emergent and often opaque behaviors, making them susceptible to cascading errors, biased decisions, and unintended interactions [9]. These risks are amplified when agents plan, coordinate, and call external tools. Beyond traditional ML concerns (safety, fairness, and interpretability), Agentic AI must contend with threats from tool integrations, prompt-level attacks, memory poisoning, impersonation, and privacy leakage in collaborative settings [10, 11]. To address these system-level risks, AI Trust, Risk, and Security Management (TRiSM) frameworks emphasize governance, explainability, security, privacy, and lifecycle controls [12, 13]. We argue that a TRiSM perspective is necessary to make AMAS deployable in sensitive domains.
Scope and Objectives
Our focus is LLM-based AMAS that demonstrate autonomous planning, tool use, memory retention, and emergent reasoning, with or without human oversight. [^4] Throughout, AMAS denotes multi-agent systems built on LLMs. We synthesize recent work on AMAS and AI TRiSM to provide a coherent, actionable reference for researchers and practitioners. Specifically, we map TRiSM pillars :Explainability, ModelOps, Security, Privacy, and their Lifecycle Governance, to AMAS workflows and failure modes. We also surface gaps where current ML governance and model-centric controls do not readily extend to multi-agent settings.
Necessity of this Survey
@@
Much of the Agentic AI literature prioritizes agent modeling, planning, and collaboration, while comparatively under-addressing adversarial robustness, lifecycle governance, decision provenance, and system-level explainability. As deployments expand in healthcare, finance, science, and public services, the absence of an integrated TRiSM perspective exposes stakeholders to opaque decision pathways and unmanaged risks. To our knowledge, few surveys examine LLM-based multi-agent systems explicitly through the lens of trust, risk, security, and governance; Table 1 situates our contribution relative to prior reviews.
@@
Table 1: Comparison with related surveys.
Survey Threats Lifecycle Gov. Explainability TRiSM Integration LLM-Specific Applications Actionable Guidance
Guo et al. (2024) [14] ~
Chen et al. (2025) [15] ~ ~
Yan et al. (2025) [16] ~
Tran et al. (2025) [17] ~
Lin et al. (2025) [18] ~ ~
Fang et al. (2025) [11] ~ ~ ~
Xi et al. (2025) [19] ~
Luo et al. (2025) [20] ~ ~ ~
Zou et al. (2025) [21] ~ ~
Wang et al. (2025) [22] ~ ~ ~ ~
@@
Contributions
We make the following contributions:
  • Unified TRiSM framework for Agentic AI. We propose and present a system-level Trust, Risk and Security Management (TRiSM) framework tailored to large-language-model-based multi-agent systems (Section 5). The framework integrates explainability, ModelOps, security, privacy and lifecycle governance, contextualising each pillar to AMAS workflows and failure modes.
  • Risk taxonomy for AMAS. We synthesise a taxonomy of threats unique to LLM-driven multi-agent systems, such as prompt injection, memory poisoning, collusive failure, emergent misbehaviour and tool-use abuse; and map these risks to corresponding controls (Section 4).
  • Evaluation template and new metrics. We propose a holistic evaluation template that covers trustworthiness, explainability, user-centred performance and inter-agent coordination. To operationalise this template, we introduce the Component Synergy Score (CSS) for measuring collaboration quality across agents and Tool Utilisation Efficacy (TUE) for assessing the correctness and efficiency of tool calls (Section 6).
  • Technique mapping and limitations. We survey and map existing explainability methods, prompt-hygiene practices, decision-provenance tools, sandboxing strategies and ModelOps pipelines to the AMAS setting. This mapping reveals where existing approaches can be adapted and where significant gaps remain (Section 8).
Building on the above contributions, we outline a research roadmap for developing scalable, verifiable and regulation-aligned agentic systems. Key directions include adversarial robustness, governance protocols, standardised benchmarks for trustworthiness and coordination, and cross-disciplinary collaboration.
2 Background, Related Work and Literature Review Methodology
2.1 AI Agents vs. Agentic AI
An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals [7], has undergone a rapid transformation in recent years. While early agents were task-specific and deterministic, modern systems have evolved substantially, increasing operational autonomy on benchmark tasks [23]. These agentic systems integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [8]. This evolution presents new opportunities but also raises challenges in coordination, control, transparency, and risk assessment across heterogeneous agents. As the shift toward agentic architectures accelerates, long-standing concerns around safety, fairness, interpretability, and systemic bias resurface [11], especially in high-stakes domains such as healthcare, law, and public services, where decisions can carry profound societal consequences.
Autonomous software agents date back to the 1990s [24]; however, the advent of LLM-driven multi-agent ecosystems since 2023 [9] has made “Agentic AI” a distinct and rapidly expanding research focus [8]. Traditional agents primarily automated limited tasks—such as information retrieval [25], data summarization [26], and dialogue response [27]—often via single-step logic [28] or scripted rules. These approaches lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem-solving.
In contrast, agentic systems coordinate collaborative agents with specialized roles (e.g., planner, coder, analyst) enabled by LLMs and tool use [29]. They dynamically decompose tasks, share context, and pursue high-level goals over long horizons [30]. This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent, decentralized behavior.
2.2 Trust, Risk, and Security Challenges in Agentic AI
The rise of Agentic AI introduces challenges that extend beyond those of traditional ML systems. Unlike deterministic agents [31], LLM-based multi-agent systems (AMAS) exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors. These characteristics necessitate robust approaches to building trust and ensuring trustworthiness in high-stakes domains such as healthcare, finance, and law. Below, we discuss several key challenges:
  1. 1.
    Trust and trustworthiness. In Agentic AI, trust refers to a user’s willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner. According to ISO [32], trustworthiness is a quality characteristic of systems and is understood as the ‘ability to meet stakeholders’ expectations in a verifiable way’. This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications. The stochastic nature of LLMs [33] further complicates trustworthiness: randomness in output generation, while sometimes promoting creativity, introduces inconsistency and limits repeatability.
  2. 2.
    Security risks in multi-agent systems. Risk surfaces expand as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [10]. For example, a multi-agent supply-chain optimizer may coordinate procurement and logistics agents yet inadvertently leak sensitive data or violate compliance protocols (e.g., [34], HIPAA [35]) if safeguards are inadequate.
  3. 3.
    Prompt injection, spoofing, and impersonation. AMAS are susceptible to security threats that exploit language-based interfaces and cooperative behaviors. One prominent threat is prompt injection [36], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior. Recent studies [37] describe “prompt infection”, in which malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [38]. Corrupted outputs can cascade, leading to data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
    Another critical vulnerability involves spoofing and impersonation [39]. Spoofing refers to falsifying identity or credentials (e.g., using a false name or address) to deceive a system, whereas impersonation involves pretending to be a specific agent or user to gain trust, access, or privileges. In AMAS, where agents rely on credentials or tokens for authentication during coordination, an adversary who steals an API key or mimics a trusted peer can issue unauthorized commands or access sensitive information.
  4. 4.
    Privacy risks in multi-agent LLM systems. LLM-based agents often process user data, proprietary business information, and other sensitive inputs to fulfill their tasks [40]. In a multi-agent context, this challenge is amplified by inter-agent information sharing (e.g., via a shared memory store or message passing), which complicates data minimization, access control, and purpose limitation.
Traditional evaluation and safety frameworks, designed for deterministic, single-agent pipelines; are inadequate for the autonomy and control properties of AMAS. To address these gaps, AI Trust, Risk, and Security Management (AI TRiSM) [12, 13] proposes lifecycle controls spanning explainability, secure model/tool orchestration, and privacy management. Rather than focusing solely on technical performance, TRiSM integrates governance practices to strengthen reliability and robustness, protect data, and align system behavior with ethical and regulatory expectations.
2.3 Survey Landscape
LLM-based multi-agent surveys (technical focus)
Prior surveys on LLM-driven AMAS predominantly address system architectures, agent capabilities, and domain-specific applications of agentic systems. For example, [14, 15] examine simulated environments, inter-agent communication protocols, and performance benchmarks. These works clarify how agents coordinate and scale but give limited treatment to adversarial threats (e.g., prompt injection, collusion), lifecycle governance (e.g., continuous monitoring, audit, regulatory compliance), and system-wide explainability, especially for emergent behaviors in complex settings.
More specialized surveys, such as [16], focus on natural-language communication, while [17] details collaboration mechanisms and practical use cases in domains like question answering, 5G networks, and Industry 5.0. Although [16] highlights security risks associated with communication channels and [17] discusses deployment contexts, both remain narrowly centered on coordination strategies. Deeper analysis is needed on robustness to adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability at scale.
Trustworthy/responsible AI surveys (broad trust focus)
Review articles addressing TRiSM-related themes (e.g., [11]; see also fairness-focused work [41]) discuss alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage). However, these surveys typically target general ML systems and overlook dynamics unique to LLM-based multi-agent settings. For instance, while they advocate privacy-preserving techniques such as differential privacy, they rarely address privacy risks stemming from inter-agent data sharing or the distributed nature of LLM-driven workflows. Moreover, critical TRiSM aspects—such as governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents), are often under-specified [42].
2.3.1 Positioning of this study
This review maps TRiSM principles to LLM-based AMAS, operationalizing governance, explainability, security, privacy, and lifecycle controls for multi-agent settings. Relative to prior surveys (Table 1), we foreground system-level risks (e.g., prompt infection, impersonation, memory poisoning), propose evaluation dimensions and composite measures for coordination and tool use, and offer actionable guidance for researchers, engineers, and policymakers.
@@ 2.4 Literature Review Methodology
We conducted a structured literature review to synthesize AI TRiSM principles for AMAS. Guided by established systematic-review practices [43, 44], this section states our research objectives and describes the data sources and search strategy, the inclusion/exclusion criteria and screening procedure, and the classification schema used for analysis.
(a) Agentic AI papers per year (2022–2025).
(b) Papers by TRiSM pillar.
(c) Topic trends in Agentic AI (agents, security, explainability).
(d) Bibliography entries per year (1987–2025).
Figure 2: Bibliographic analysis of Agentic AI research. (a) Annual trend in Agentic AI publications, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics. (c) Topic-specific trends in agents, security, and explainability. (d) Overall growth in bibliographic entries, highlighting acceleration since 2020.
@@
  • RQ1: What trust, risk, and security challenges are specific to LLM-based AMAS?
@@
We systematically searched major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search covered publications from January 2022 to May 2025, with select seminal works prior to 2022 included for foundational context. We used Boolean queries and expanded keywords to capture the intersection of LLM-based multi-agent systems and TRiSM considerations. Example query template:
("Agentic AI" OR "multi-agent systems" OR "multi-agent LLMs" OR "AI agents" OR "autonomous agents" OR "intelligent agents" OR "collaborative AI" OR "distributed AI" OR "LLM-based agents" OR "agent coordination" OR "agent-based modeling" OR "swarm intelligence")
@@
("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR "governance" OR "oversight" OR "compliance" OR "explainability" OR "interpretability" OR "transparency" OR "privacy" OR "data protection" OR "robustness" OR "accountability" OR "ethical AI" OR "fairness" OR "adversarial attacks" OR "regulatory compliance" OR "bias mitigation" OR "system reliability")
@@
Two-stage screening (title/abstract, then full-text) used these criteria. Inclusion:
  • Studies published 2022–2025 (capturing the post-2022 LLM wave), plus select pre-2022 foundational work.
  • Studies explicitly discussing Agentic AI, multi-agent, or LLM-powered systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
  • Peer-reviewed articles and widely cited preprints or reports from credible sources (e.g., NeurIPS, IEEE, ACM, Nature, governmental/standards bodies, arXiv).
Exclusion:
  • Papers focused solely on traditional, rule-based, or symbolic agents without LLM integration or emergent coordination.
  • Studies addressing only low-level ML components (e.g., training optimization, architecture design) with no agentic behavior or TRiSM concerns.
  • Non-English papers or those lacking sufficient metadata (e.g., missing abstract or publication details).
For empirical papers, we assessed clarity of threat models, reproducibility (code/data), evaluation rigor (baselines, ablations), and external validity. For surveys, we assessed coverage breadth, structure, and treatment of risks/governance. After screening titles and abstracts, we shortlisted 250 papers; full-text review yielded 180 primary studies addressing one or more TRiSM pillars. We also incorporated relevant technical reports and authoritative web resources to contextualize the rapidly evolving Agentic-AI/TRiSM landscape. A bibliometric summary appears in Figure 2.
@@
To ensure reliability, we adapted quality criteria from established guidelines [45]. Each paper was evaluated on: (1) clear objectives for agentic or LLM-based systems; (2) well-documented methods/experiments or architecture (reproducibility where applicable); (3) substantive engagement with at least one of trust, risk, security, explainability, or privacy; (4) empirical, theoretical, or normative contributions relevant to Agentic-AI governance. Each criterion was rated low, medium, or high; studies rated low on more than two criteria were excluded or flagged as contextual only. Quality ratings were assigned independently by four reviewers; discrepancies were resolved via discussion (with inter-rater agreement checked on a subset).

### v5

@@ Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around key pillars: Explainability, ModelOps, Security, Privacy and their Lifecycle Governance, each contextualized to the challenges of AMAS. A risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To make coordination and tool use measurable in practice, we propose two metrics: the Component Synergy Score (CSS), which captures inter-agent enablement, and the Tool Utilization Efficacy (TUE), which evaluates whether tools are invoked correctly and efficiently. We further discuss strategies for improving explainability in Agentic AI, as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, highlighting key directions to align emerging systems with TRiSM principles-ensuring safety, transparency, and accountability in their operation.
@@ 1 Introduction
AI governance has moved from aspiration to obligation. The EU Artificial Intelligence Act [EU_AI_Act2025] entered into force on 1 August 2024 and applies progressively (e.g., prohibited AI practices and AI literacy obligations from 2 February 2025, and governance rules plus obligations for general-purpose AI models from 2 August 2025) [^3] EU AI Act; meanwhile, organizations are adopting management frameworks such as ISO/IEC 42001:2023 [iso42001] and the NIST AI Risk Management Framework (AI RMF 1.0) [ai2023artificial] to operationalize risk controls. In parallel, enterprise use of AI continues to expand, for example, recent global surveys [McKinsey_StateOfAI2025, Stanford_HAI_AIIndex2025, Gartner_AI_TRiSM2024] report that a substantial majority of organizations use AI in at least one business function. Critically, the emerging paradigm of AI agents, which are systems that autonomously plan, reason, and execute complex tasks, introduces qualitatively different risks than traditional AI applications. These trends demand system-level trust, risk, and security management tailored to agentic systems.
Figure 1: Agentic AI research growth on arXiv (2019–2024), showing multi-agent systems and LLM-based agent publications with marked acceleration following ChatGPT’s release in November 2022.
Figure 2: Google search interest for “AI Agents” vs “Agentic AI” (2022–2024), illustrating rising public awareness with key milestones including ChatGPT launch and Gartner’s recognition of Agentic AI as the top technology trend for 2025.
@@
An AI agent is a computational entity that perceives its environment and acts to achieve goals [russell1997rationality]. What were once task-specific and largely deterministic programs have rapidly evolved into LLM-powered agentic systems with planning, tool use, and persistent memory [sapkota2025ai]. In this paper, we refer to such LLM-based, coordinating agent systems as Agentic Multi-Agent Systems (AMAS). These systems orchestrate multiple specialized agents to address complex tasks, coordinate functions and roles, and adapt workflows through interactions with tools, users, and other agents.
TRiSM for Agentic AI
@@
AMAS exhibit emergent and often opaque behaviors, making them susceptible to cascading errors, biased decisions, and unintended interactions [yang2024multi]. These risks are amplified when agents plan, coordinate, and call external tools. Beyond traditional ML concerns (safety, fairness, and interpretability), Agentic AI must contend with threats from tool integrations, prompt-level attacks, memory poisoning, impersonation, and privacy leakage in collaborative settings [lim2012memory, fang2025trustworthyaisafetybias]. To address these multifaceted system-level risks, AI Trust, Risk, and Security Management (TRiSM) frameworks provide integrated approaches emphasizing governance, explainability, security, privacy, and lifecycle controls [IBM2025AITRiSM, habbal2024artificial].
Integrating TRiSM principles into the design and deployment of Agentic AI systems is increasingly urgent. Figure 1 demonstrates explosive growth in arXiv publications, with multi-agent systems papers increasing from 890 in 2019 to 18,500 in 2024, while LLM-based agent papers surged from near zero to 9,800 following ChatGPT’s release. This academic interest is mirrored by public awareness, as shown in Figure 2, where Google search trends for both “AI Agents” and “Agentic AI” reached peak levels by late 2024. Gartner’s designation of Agentic AI as a top strategic technology trend for 2025, alongside AI governance platforms within its TRiSM framing [^4] Gartner Market Guide for AI TRiSM, and its 2026 emphasis on multiagent systems, digital provenance, and AI security platforms  [^5] Gartner Top Technology Trends 2026 reflect a growing consensus: as AI systems gain autonomy, robust trust, risk mitigation, and security controls become indispensable. We emphasize that a TRiSM perspective is essential to make AMAS deployable in sensitive domains.
1.1 Necessity and Scope of this Survey
@@
Much of the Agentic AI literature prioritizes agent modeling, planning, and collaboration, while comparatively less attention is given to adversarial robustness, lifecycle governance, decision provenance, and system-level explainability. This survey addresses this gap by focusing on AMAS [^6] Throughout, AMAS denotes multi-agent systems built on LLMs. that demonstrate autonomous planning, tool use, memory retention, and emergent reasoning. As deployments expand in healthcare, finance, science, and public services, the absence of an integrated TRiSM perspective exposes stakeholders to opaque decision pathways and unmanaged risks. To our knowledge, no prior survey examines LLM-based multi-agent systems explicitly through the lens of trust, risk, security, and governance. Table 1 situates our contribution relative to prior work.
1.2 Contributions
The contributions of this study are manifold:
  • Unified TRiSM framework for Agentic AI. We present a system-level Trust, Risk and Security Management (TRiSM) framework tailored to AMAS (Section 4). The framework integrates explainability, ModelOps, security, privacy and lifecycle governance, contextualizing each pillar to AMAS workflows and failure modes.
  • Risk taxonomy for AMAS. We synthesize a taxonomy of threats unique to AMAS, including prompt injection, memory poisoning, collusive failure, emergent misbehavior and tool-use abuse, and map these risks to corresponding controls (Section 3).
  • Evaluation template and new metrics. We propose a holistic evaluation template covering trustworthiness, explainability, user-centered performance and inter-agent coordination. To operationalize this template, we introduce the Component Synergy Score (CSS) for measuring collaboration quality across agents and Tool Utilization Efficacy (TUE) for assessing the correctness and efficiency of tool calls (Section 5).
  • Technique mapping and gap analysis. We survey and map existing explainability methods, prompt-hygiene practices, decision-provenance tools, sandboxing strategies and ModelOps pipelines to the AMAS setting, revealing where existing approaches can be adapted and where significant gaps remain (Section 7).
Building on these contributions, we outline a research roadmap for developing scalable, verifiable and regulation-aligned agentic systems. Key directions include adversarial robustness, governance protocols, standardized benchmarks for trustworthiness and coordination, and cross-disciplinary collaboration.
1.3 Related Work
Prior surveys on LLM-based multi-agent systems mainly emphasize architectures, agent capabilities, communication, and domain applications [guo2024largelanguagemodelbased, chen2025surveyllmbasedmultiagentsystem, yan2025selftalkcommunicationcentricsurveyllmbased, tran2025multiagentcollaborationmechanismssurvey]. In parallel, broader trustworthy/responsible AI surveys discuss themes aligned with TRiSM, such as alignment, fairness, and privacy defenses; but typically focus on general ML systems and do not model AMAS-specific dynamics (e.g., inter-agent information sharing, tool-mediated actions, and emergent coordination) [fang2025trustworthyaisafetybias, raza2024exploring]. As a result, key gaps remain in adversarial robustness, governance and accountability for multi-agent deployments, and explainability at scale. In comparison to these works (Table 1), our study addresses these gaps by mapping TRiSM principles to AMAS and operationalizing governance, explainability, security, privacy, and lifecycle controls for multi-agent settings.
@@ 1.3 Related Work
Table 1: Comparison with representative related surveys on LLM-based agentic systems.
Survey Threats Lifecycle Explain. TRiSM LLM-Spec. Apps. Actionable
Guo et al. (2024) [guo2024largelanguagemodelbased] ~ ~
Chen et al. (2024) [chen2025surveyllmbasedmultiagentsystem] ~ ~
Yan et al. (2025) [yan2025selftalkcommunicationcentricsurveyllmbased] ~ ~
Tran et al. (2025) [tran2025multiagentcollaborationmechanismssurvey] ~
Lin et al. (2025) [lin2025creativityllmbasedmultiagentsystems] ~ ~
Fang et al. (2025) [fang2025trustworthyaisafetybias] ~ ~ ~
Xi et al. (2025) [xi2025rise] ~ ~
Luo et al. (2025) [luo2025large] ~ ~ ~
Zou et al. (2025) [zou2025llm] ~ ~
Wang et al. (2025) [wang2025surveyllmbasedagentsmedicine] ~ ~
Wang et al. (2025) [11081880] ~
Yu et al. (2025) [yu2025survey] ~
@@ 1.4 Literature Review Methodology
We conducted a structured literature review to synthesize AI TRiSM principles for AMAS. Following established systematic review guidance [boland2017doing, keele2007guidelines], we describe our data sources, search strategy, screening criteria, and classification schema.
@@
  • RQ1: What trust, risk, and security challenges are specific to AMAS?
@@
We systematically searched major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search covered publications from January 2022 to December 2025, with influential foundational works prior to 2022 included for theoretical context. We used a two-block Boolean query: (agentic/AMAS terms) AND (TRiSM terms).
("agentic AI" OR "multi-agent systems" OR "multi-agent LLMs" OR "AI agents"
OR  "autonomous agents" OR "intelligent agents" OR "collaborative AI" OR
 "distributed AI" OR "LLM-based agents" OR "agent coordination" OR
 "agent-based modeling" OR "swarm intelligence")
@@
("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR
 "governance" OR "oversight" OR "compliance" OR "explainability" OR
 "interpretability" OR "transparency" OR "privacy" OR "data protection" OR
 "robustness" OR "accountability" OR "ethical AI" OR "fairness" OR
 "adversarial attacks" OR "regulatory compliance" OR "bias mitigation" OR
 "system reliability")
@@
We applied a two-stage screening process (title/abstract, then full text).
Inclusion criteria:
  • Published between 2022–2025 (plus influential foundational papers published before 2022).
  • Focused on agentic AI, multi-agent, or LLM-powered systems and addressed at least one TRiSM aspect (trust, risk, security, governance, explainability, or privacy).
  • Peer-reviewed papers, widely cited preprints, technical reports from reputable venues or organizations (e.g., IEEE/ACM conferences and journals, standards bodies, and government sources).
Exclusion criteria:
  • Traditional rule-based or symbolic agents without LLM-based autonomy or coordination.
  • Papers limited to low-level model training/architecture details with no agentic behavior or TRiSM relevance.
  • Non-English papers or items with insufficient bibliographic information (e.g., missing abstract or venue details).
For empirical studies, we verified that the threat model and evaluation were clearly described and assessed reproducibility when code or data were available. For survey papers, we evaluated coverage, structure, and treatment of risks and governance. After title/abstract screening, we shortlisted 250 papers; full-text screening yielded 180 primary studies covering one or more TRiSM pillars, supplemented by relevant standards and technical reports for regulatory context.
@@
To ensure reliability, we adapted quality criteria from established guidelines [kitchenham2004procedures]. Each paper was evaluated on: (1) clear objectives for AMAS; (2) well-documented methods/experiments or architecture (reproducibility where applicable); (3) substantive engagement with at least one of trust, risk, security, explainability, or privacy; (4) empirical, theoretical, or normative contributions relevant to agentic AI governance. Each criterion was rated low, medium, or high; studies rated low on more than two criteria were excluded or flagged as contextual only. Quality ratings were assigned independently by four reviewers; discrepancies were resolved via discussion.
Bibliometric Overview
The results of our screening process reveal the rapid acceleration of agentic AI research. The publication landscape (Figure 3) shows journal articles dominating at 61.2%, with leading venues and publishers detailed in Figure 4. Key research themes, including large language models, multi-agent systems, and security, are visualized in Figure 5a, while Figure 6 confirms the unprecedented growth of research activity in recent years.
(a) Publication trends (1985–2025)
(b) Publication type distribution
Figure 3: Publication landscape: (a) Long-term trends showing accelerating research output over four decades; (b) Distribution of publication types, with journal articles dominating at 61.2%.
(a) Top publication venues
(b) Publisher distribution
Figure 4: Publication outlets: (a) Top venues including NeurIPS, ACM Computing Surveys, and AAAI; (b) Publisher distribution led by Elsevier, Springer, and IEEE.
(a) Research topics and keywords
(b) Publication type evolution over time
Figure 5: Research themes and evolution: (a) Word cloud highlighting key topics including large language models, multi-agent systems, security, and explainability; (b) Temporal distribution of publication types.
Figure 6: Comparison of recent publications (2020–2025) with historical distribution by period, demonstrating the rapid acceleration of agentic AI research in recent years.
Figures 7 and 8 provide a roadmap for the paper, which progresses from architectural foundations to risk analysis, evaluation, and governance.
TRiSM for Agentic AIFundamentalsThreats & RisksTRiSM FrameworkEvaluation
$\circ$ AMAS Architecture
$\circ$ LLM Core
$\circ$ Planning/Reasoning
$\circ$ Memory Module
$\circ$ Tool Interface
$\circ$ Adversarial Attacks
$\circ$ Data Leakage
$\circ$ Agent Collusion
$\circ$ Emergent Behavior
$\circ$ Memory Poisoning
$\circ$ Explainability
$\circ$ ModelOps
$\circ$ Application Security
$\circ$ Model Privacy
$\circ$ Governance
$\circ$ Trustworthiness
$\circ$ CSS Metric
$\circ$ TUE Metric
$\circ$ Compliance
$\circ$ Auditability
Figure 7: Taxonomy of TRiSM for LLM-based Agentic Multi-Agent Systems.
Figure 8: TRiSM taxonomy for Agentic AI (AMAS) presented in this review. The paper progresses through six phases: (i) Fundamentals of Agentic AI: distinguishing AMAS from traditional agents and establishing LLM-based architectures (Section 2); (ii) AI Trust, Risk, and Security Management (TRiSM) framework pillars: Explainability, ModelOps, Application Security, Model Privacy, and Lifecycle Governance, introduced conceptually (Section 4); (iii) Trust management via explainability techniques (Section 4.2.1); (iv) Threats and risks taxonomy with real-world case studies (Section 3); (v) Evaluation metrics encompassing trustworthiness, coordination, and composite measures (Section 5); (vi) Technical implementation of security/privacy mechanisms (Section 6); and (vii) Compliance and governance integration (Section 7).